OpenStudy (anonymous):

Calculate y'. xy^4+x^2y=x+3y

OpenStudy (earthcitizen):

have you heard of implicit differentiation ?

OpenStudy (anonymous):

yes but i dont get any of the answer choices as i answer it

mathslover (mathslover):

\[\Huge{xy^4+x^2y=x+3y}\] \[\Huge{xy(y^3+x)=x+3y}\] \[\Huge{xy(y^3+x)-x=3y}\] \[\Huge{x(y(y^3+x)-x)=3y}\] \[\Huge{\frac{x^2y}{3}(y^4+xy-1)=y}\]

OpenStudy (anonymous):

see that isnt one of my answer choices though

OpenStudy (blockcolder):

\[\large4xy^3y'+y^4+2xy+x^2y'=1+3y'\\ \large 4xy^3y'+x^2y'-3y'=1-2xy-y^4\\ \large y'(4xy^3+x^2-3)=1-2xy-y^4\\ \Large y'=\frac{1-2xy-y^4}{4xy^3+x^2-3}\] This isn't one of the choices?